OpenStudy (anonymous):

I need help factoring the polynomial completely: 12(x-3)^2-2(x-3)-30

OpenStudy (asnaseer):

let y=x-3, so you then get 12y^2-2y-30=2(2y+3)(3y-5) replace y to get expression in x.

OpenStudy (anonymous):

let x-3 = u 12u^2 - 2u - 30 = 2(6u^2 - u - 15) = 2(3u - 5 )(2u + 3 ) now replace u with x-3 2(3x- 9 - 5)(2x - 6 +3) = 2(3x - 14)(2x - 3)

OpenStudy (anonymous):

Thanks so much, that helped me a ton!

OpenStudy (anonymous):

yw